# and Inverse in Commutative Analogues of Clifford Algebras

### v2

@@ 2.1.1. Grades and projections
We define projection maps .${\langle{.}\rangle}_{k}:{{{K}}_{{p},{q}}}\to{{{K}}^{k}_{{p},{q}}}$; $U\mapsto{\langle{U}\rangle}_{k}$ which are linear maps which return the grade $k$ part of $U$. In general,
@@ 2.2. Tensor product decomposition
An described in [36], an immediate consequence of commutativity is that one can decompose any ${{{K}}_{{p},{q}}}$ as tensor products of ‘$p$${{{K}}_{{1},{0}}}$s and ‘$q$${{{K}}_{{0},{1}}}$s.
@@ 2.3. Operations of conjugation
We define $n=p+q$ operations of conjugation in ${{{K}}_{{p},{q}}}$ where each operation of conjugation negates one generator. More explicitly, for $l\in\{1,2,\ldots,{n}\}$, we define (.${{(.)}^{(l)}}:{{{K}}_{{p},{q}}}\to{{{K}}_{{p},{q}}}$ such that ${{U}^{(l)}}=U|_{e_{l}\to-e_{l}}$.
@@
Note that the set of all superpositions of operations of conjugates form an Abelian group (as mentioned in [4, 11] for the particular case of multicomplex spaces). The group operation in the group is superposition of operations of conjugation and the identity element of this group is the operation of conjugation ${{(.)}^{(\{\})}}$.
Also, we make an observation. The only operation of conjugation that can eliminate all terms involving the generator $e_{k}$ is (.${{(.)}^{(k)}}$. We usually eliminate $e_{k}$ in an element $U\in{{{K}}_{{p},{q}}}$ by multiplying U with the conjugate ${{U}^{(k)}}$ as presented in the following lemma.
@@ 3.2.1. Tensor product of linear operators
Let $T_{1}:V_{1}\to W_{1}$ and $T_{2}:V_{2}\to W_{2}$ be two linear operators. The tensor product of $T_{1}$ and $T_{2}$ is the linear map $T_{1}\otimes T_{2}:V_{1}\otimes V_{2}\to W_{1}\otimes W_{2}$ such that
@@
If we fix basis for $V_{1},W_{1},V_{2},W_{2}$, we can associate a matrix to $T_{1}\otimes T_{2}$. Let $\mathrm{A}$, $\mathrm{B}$ be the matrices corresponding to $T_{1}$ and $T_{2}$ respectively. Then the matrix corresponding to $T_{1}\otimes T_{2}$ is the Kronecker product of $\mathrm{A}$ and $\mathrm{B}$, i.e., $\mathrm{A}\otimes\mathrm{B}$. It is the block matrix:
@@ 3.3. Representation of ${{{K}}_{{p},{q}}}$
where each $e_{j}$ corresponds to the Kronecker product of ‘$n$$I$s except with $\mathrm{J}_{+}$ at $j^{th}$ spot if $e_{j}^{2}=+1$ or $\mathrm{J}_{-}$ at $j^{th}$ spot if $e_{j}^{2}=-1$.
@@ 5.1. Formulas for trace
This is because the only basis vector having non-zero trace is $1$. All other basis vectors have trace 00.
@@ 5.2.1. Inspiration from examples with small $n$
Finally, $\rm Det(U)=\rm det(\beta(U))=((a+{a_{1}})^{2}+({a_{2}}+a_{12})^{2})((a-{a_{1}})^{2}+({a_{2}}-a_{12})^{2})$ $=U{{U}^{(1)}}{{U}^{(2)}}{{U}^{(1)(2)}}$. The right-hand side of this expression has already been used to determine the norm of commutative quaternions in [20, 5].
@@ 5.2.2. The formula for determinant
In the above formula, the right-hand side is the product of all conjugates of $U$. In total, there are $N=2^{n}$ productants in right-hand side corresponding to total number of different superpositions of $n$ operations of conjugation defined in Section 2.3. Using the notation introduced in (2.7), we can express the formula for determinant as
@@ 5.3.1. Explicit formulas for coefficients of characteristic polynomials
Let $\psi_{U}(\lambda)={\sum_{k=0}^{N}}{{c_{k}(U)}{\lambda^{k}}}$ be the characteristic polynomial where $c_{k}(U)$s the coefficients of characteristic polynomial as defined in Section 4. We present explicit formulas for them as the following theorem.
@@
We have
@@
where the sum is over all distinct subsets of $\{1,2,\ldots,{n}\}$ $N-k$ at a time.
@@
Notice that $c_{0}(U)$ gives us the formula for determinant of $U$ that we know from Section 5.2.2 and $-c_{N-1}(U)$ is the trace of $U$ and thus gives us the formula for trace of $U$ which we had got in Theorem 5.2.
@@
Note that we could get the same formulas for coefficients of characteristic polynomial (5.7) by using Vieta’s formulas [^4] Compare with the more non-trivial non-commutative Vieta formulas for the case of Clifford algebras presented in the paper [38].. Since all conjugates of $U$, $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ satisfy the characteristic polynomial of $U$, $\psi_{U}$ one can argue using Vieta’s Theorem that the coefficients of characteristic polynomial $\psi_{U}$ are symmetric polynomials of $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ i.e., the coefficients of characteristic polynomial do not change under a permutation of the elements in the set $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ which are precisely the expressions (5.7).
@@ 6. Multiplicative inverses in ${{{K}}_{{p},{q}}}$
Define $F:{{{K}}_{{p},{q}}}\to{{{K}}_{{p},{q}}}$; $F(U):={{U}^{(1)}}{{U}^{(2)}}\cdots{{U}^{(1)(2)}}\cdots{{U}^{(1)(2)(3)\cdots(n)}}$ such that $\rm Det(U)$ $=UF(U)$. We call $F(U)$ the adjoint of the element $U$.
@@ 6.1. Explicit formulas for multiplicative inverses
The proof follows from that fact that $\rm Det(U)=UF(U)$. Therefore, if $\rm Det(U)\neq 0$, then $U^{-1}=\frac{F(U)}{\rm Det(U)}.$
We give explicit formulas for $F(U)$ for $n\leq 4$ below:
$F(U)=\left\{\begin{array}[]{ll}{{U}^{(1)}},&\mbox{if n = 1};\\ {{U}^{(1)}}{{U}^{(2)}}{{U}^{(12)}},&\mbox{if n = 2};\\ {{U}^{(1)}}{{U}^{(2)}}{{U}^{(3)}}{{U}^{(12)}}{{U}^{(23)}}{{U}^{(13)}}{{U}^{(123)}},&\mbox{if n = 3};\\ {{U}^{(1)}}{{U}^{(2)}}{{U}^{(3)}}{{U}^{(4)}}{{U}^{(12)}}{{U}^{(13)}}{{U}^{(14)}}{{U}^{(23)}}\\ \,\,\cdot\,{{U}^{(24)}}{{U}^{(34)}}{{U}^{(123)}}{{U}^{(134)}}{{U}^{(124)}}{{U}^{(234)}}{{U}^{(1234)}},&\mbox{if n = 4}.\end{array}\right.$
@@
This follows from $F({{U}^{(A)}})={{F(U)}^{(A)}}$. ∎
@@ 7. Conclusions
The presented explicit formulas for determinant in ${{{K}}_{{p},{q}}}$ (formula (5.3)) are a single product of conjugates. This is in contrast with the formulas for determinant in Clifford algebras ${{{C}\!{\ell}}_{{p},{q}}}$ where explicit formulas for determinant are in general a linear combination of terms involving conjugates (we have an algorithm to calculate it, check Theorem 4 in [37] and Theorem 3 in [38]). For $p+q\leq 5$, formula for determinant is a single product (check [37]), for $p+q=6$ the formula is a linear combination of 2 terms (check [1]) and for $p+q\geq 7$ the number of terms in the formula for determinant is not known. Our main aim for this paper was to address the multiplicative inverse problem in ${{{K}}_{{p},{q}}}$ i.e. how to find multiplicative inverses in ${{{K}}_{{p},{q}}}$ because we believe that solving the multiplicative inverse problem in commutative analogues might help us better understand the multiplicative inverse problem in ${{{C}\!{\ell}}_{{p},{q}}}$ and allow us to find more explicit formulas for determinants for $p+q\geq 7$ with least number of terms in the expression for determinant. The connection between commutative analogues of Clifford algebras and Clifford algebras is made by some subalgebras of ${{{C}\!{\ell}}_{{p},{q}}}$ being isomorphic to ${{{K}}_{{p^{\prime}},{q^{\prime}}}}$ that cause trouble in Clifford algebras by forcing the formula for determinant to be a linear combination rather than a single product (check out [1]). This is a topic to be further explored.

### v3

@@ 2.1.1. Grades and projections
We define projection maps ${\langle{.}\rangle}_{k}:{{{K}}_{{p},{q}}}\to{{{K}}^{k}_{{p},{q}}}$; $U\mapsto{\langle{U}\rangle}_{k}$ which are linear maps which return the grade $k$ part of $U$. In general,
@@ 2.2. Tensor product decomposition
[…]
@@ 2.3. Operations of conjugation
We define $n=p+q$ operations of conjugation in ${{K}}_{{p},{q}}$} where each operation of conjugation negates one generator. More explicitly, for $l\in\{1,2,\ldots,{n}\}$, we define ${{(.)}^{(l)}}:{{{K}}_{{p},{q}}}\to{{{K}}_{{p},{q}}}$ such that ${{U}^{(l)}}=U|_{e_{l}\to-e_{l}}$.
@@
[…]
Also, we make an observation. The only operation of conjugation that can eliminate all terms involving the generator $e_{k}$ is ${(.)}^{(k)}$}. We usually eliminate $e_{k}$ in an element $U\in{{{K}}_{{p},{q}}}$ by multiplying U with the conjugate ${U}^{(k)}$} as presented in the following lemma.
@@ 3.1. Representations of ${{K}}_{{1},{0}}$} and ${{K}}_{{0},{1}}$}
Remark 3.1.
Note that in this paper we consider only real representations of ${{K}}_{{p},{q}}$}, since we want to deal with a real determinant. For example, ${{K}}_{{0},{1}}$$\simeq{\mathbb{C}}$ also has a complex representation $1\mapsto 1,e_{1}\mapsto i$. Another example is ${{K}}_{{0},{2}}$}, the bicomplex space, whose complex representation is discussed in Section 2.4 of [21]. Computing determinant from these complex representations would give a complex determinant.
@@ 3.2.1. Tensor product of linear operators
Let $T_{1}:V_{1}\to W_{1}$ and $T_{2}:V_{2}\to W_{2}$ be two linear operators. The tensor product of $T_{1}$ and $T_{2}$ is the linear map $T_{1}\otimes T_{2}:V_{1}\otimes V_{2}\to W_{1}\otimes W_{2}$ such that [^2] Note that a general element of $V_{1}\otimes V_{2}$ is of the form $\sum_{i,j}{a_{ij}({v_{i}}\otimes{v_{j}}})$. Since the map $T_{1}\otimes T_{2}$ is linear, defining its action on ‘pure tensors’ $v_{i}\otimes v_{j}$ suffices.
@@
If we fix basis for $V_{1},W_{1},V_{2},W_{2}$, we can associate a matrix to $T_{1}\otimes T_{2}$. Let $\mathrm{A}$, $\mathrm{B}$ be the matrices corresponding to $T_{1}$ and $T_{2}$ respectively. Then the matrix corresponding to $T_{1}\otimes T_{2}$ is the Kronecker product of $\mathrm{A}$ and $\mathrm{B}$, i.e., $\mathrm{A}\otimes\mathrm{B}$. It is the block matrix [^3] Note that sometimes the Kronecker product of two matrices $A$ and $B$ is defined to be the block matrix with blocks $a_{ij}B$ instead. This does not affect the results presented in this paper because one can reverse the ordering of the generators $\{e_{1},e_{2},\ldots,e_{n}\}$ to $\{e_{n},e_{n-1},\ldots,e_{1}\}$ which does not change the generated algebra ${{K}}_{{p},{q}}$} but the Kronecker products, as defined in this paper, in the tensor product representation will then match with the Kronecker products defined in the alternate way.:
@@ 3.3. Representation of ${{K}}_{{p},{q}}$}
where each $e_{j}$ corresponds to the Kronecker product of ‘nnIIs except with $\mathrm{J}_{+}$ at $j^{th}$ spot if $e_{j}^{2}=+1$ or $\mathrm{J}_{-}$ at $j^{th}$ spot if $e_{j}^{2}=-1$.
@@ 5.1. Formulas for trace
This is because the only basis vector having non-zero trace is $1$. All other basis vectors have trace 0.
@@ 5.2.1. Inspiration from examples with small $n$
Finally, ${\rm Det}(U)={\rm det}(\beta(U))=((a+{a_{1}})^{2}+({a_{2}}+a_{12})^{2})((a-{a_{1}})^{2}+({a_{2}}-a_{12})^{2})$ $U{{U}^{(1)}}{{U}^{(2)}}{{U}^{(1)(2)}}$. The right-hand side of this expression has already been used to determine the norm of commutative quaternions in [20, 5].
@@ 5.2.2. The formula for determinant
In the above formula, the right-hand side is the product of all conjugates of $U$. In total, there are $N=2^{n}$ factors in right-hand side corresponding to total number of different superpositions of $n$ operations of conjugation defined in Section 2.3. Using the notation introduced in (2.7), we can express the formula for determinant as
@@ 5.3.1. Explicit formulas for coefficients of characteristic polynomials
Let $\psi_{U}(\lambda)={\sum_{k=0}^{N}}{{c_{k}(U)}{\lambda^{k}}}$ be the characteristic polynomial where $c_{k}(U)$s are the coefficients of the characteristic polynomial as defined in Section 4. We present explicit formulas for them as the following theorem.
@@
Let $\psi_{U}(\lambda)={\sum_{k=0}^{N}}{{c_{k}(U)}{\lambda^{k}}}$ be the characteristic polynomial of $U\in{{{K}}_{{p},{q}}}$. Then
@@
where the sum is over all distinct subsets of $\{1,2,\ldots,{n}\}$ NkN-k at a time.
@@
Notice that $c_{0}(U)$ gives us the formula for determinant of $U$ that we know from Section 5.2.2 and $c_{N-1}(U)$ is the trace of $U$ and thus gives us the formula for trace of $U$ which we had got in Theorem 5.2.
@@
Note that we could get the same formulas for coefficients of characteristic polynomial (5.7) by using Vieta’s formulas [^4] Compare with the more non-trivial non-commutative Vieta formulas for the case of Clifford algebras presented in the paper [38].. Since all conjugates of $U$, $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ satisfy the characteristic polynomial of $U$, $\psi_{U}$ one can argue using Vieta’s Theorem that the coefficients of characteristic polynomial $\psi_{U}$ are symmetric polynomials of $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ i.e., the coefficients of the characteristic polynomial do not change under a permutation of the elements in the set $\{{{U}^{(A)}}\;|\;A\subseteq\{1,2,\ldots,{n}\}\}$ which are precisely the expressions (5.7).
@@ 6. Multiplicative inverses in ${{K}}_{{p},{q}}$}
Define ${\rm Adj}:{{{K}}_{{p},{q}}}\to{{{K}}_{{p},{q}}}$,
${\rm Adj}(U):={{U}^{(1)}}{{U}^{(2)}}\cdots{{U}^{(1)(2)}}\cdots{{U}^{(1)(2)(3)\cdots(n)}},$ (6.1)
such that ${\rm Det}(U)$ $U{\rm Adj}(U)$. We call ${\rm Adj}(U)$ the adjoint of the element $U$.
@@ 6.1. Explicit formulas for multiplicative inverses
The proof follows from that fact that ${\rm Det}(U)=U{\rm Adj}(U)$. Therefore, if ${\rm Det}(U)\neq 0$, then $U^{-1}=\frac{{\rm Adj}(U)}{{\rm Det}(U)}.$
We give explicit formulas for ${\rm Adj}(U)$ for $n\leq 4$ below:
[…]
@@
This follows from ${\rm Adj}({{U}^{(A)}})={{{\rm Adj}(U)}^{(A)}}$. ∎
@@ 6.2. A classification of elements in ${{K}}_{{p},{q}}$} based on $\mathrm{Det}$
Remark 6.4.
Note that we can also calculate all coefficients of the characteristic polynomial, adjoint, and inverse of $U\in{{{K}}_{{p},{q}}}$ using the well-known Faddeev–LeVerrier algorithm (a recursive method to calculate the coefficients of the characteristic polynomial from the matrix theory) [^5] Similarly to how we do it in (noncommutative) Clifford algebras, see [37].:
$U_{(1)}:=U,\quad U_{(k+1)}:=U(U_{(k)}-c_{k}),\quad c_{k}=\frac{N}{k}\langle U_{(k)}\rangle_{0},\quad 1\leq k\leq N,$ (6.2)
in particular,
${\rm Det}(U)=-U_{(N)}=-c_{N}=U(c_{N-1}-U_{(N-1)}).$ (6.3)
If ${\rm Det}(U)\neq 0$, then
$U^{-1}=\frac{{\rm Adj}(U)}{{\rm Det}(U)},\qquad{\rm Adj}(U)=c_{N-1}-U_{(N-1)}.$ (6.4)
@@ 7. Conclusions
The presented explicit formulas for determinant in ${{K}}_{{p},{q}}$} (5.3) and (5.4) are a single product of conjugates. This is in contrast with the formulas for determinant in Clifford algebras ${{C}\!{\ell}}_{{p},{q}}$} where explicit formulas for determinant are in general a linear combination of terms involving conjugates (we have an algorithm to calculate it, check Theorem 4 in [37] and Theorem 3 in [38]). For $p+q\leq 5$, formula for determinant is a single product (check [37]), for $p+q=6$ the formula is a linear combination of 2 terms (check [1]) and for $p+q\geq 7$ the number of terms in the formula for determinant is not known. Our main aim for this paper was to address the multiplicative inverse problem in ${{K}}_{{p},{q}}$} i.e. how to find multiplicative inverses in ${{K}}_{{p},{q}}$} because we believe that solving the multiplicative inverse problem in commutative analogues might help us better understand the multiplicative inverse problem in ${{C}\!{\ell}}_{{p},{q}}$} and allow us to find more explicit formulas for determinants for $p+q\geq 7$ with least number of terms in the expression for determinant. The connection between commutative analogues of Clifford algebras and Clifford algebras is made by some subalgebras of ${{C}\!{\ell}}_{{p},{q}}$} being isomorphic to ${{K}}_{{p^{\prime}},{q^{\prime}}}$} that cause trouble in Clifford algebras by forcing the formula for determinant to be a linear combination rather than a single product (check out [1]). This is a topic to be further explored.
@@ Acknowledgment
The authors are grateful to the anonymous reviewers for their careful reading of the paper and helpful comments on how to improve the presentation.